Problem 5-13

A construction project is broken down into the following 10 activities:

ACTIVITY IMMEDIATE
PREDECESSOR
TIME (WEEKS)
1 5
2 1 3
3 1 4
4 1 2
5 2,3 6
6 3 4
7 4 2
8 5 4
9 6,7 7
10 8,9 5

b-1. What is the critical path?

1-3-5-8-10
1-3-6-9-10
1-4-7-9-10
1-2-5-8-10

b-2. How many weeks will the project take to complete?

Number of weeks            

c. If activities 1 and 10 cannot be shortened, but activities 2 through 9 can be shortened to a minimum of one week each at a cost of $10,000 per week, which activities would you shorten to cut the project by four weeks? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer.)

Cut activity 3 by 3 weeks.
and then reduce activity 6 or 9 by one week.
Cut activity 6 by 4 weeks.
Reduce activities 6 and 9 by a total of 3 weeks.
and then reduce activity 3 by one week.
Answer #1

3 N 5 4 5 8 5 5 10 4 1 4 3 6 7 9 2 2 4 7

Answer b-1=

Different paths with durations =

1-2-5-8-10=5+3+6+4+5=23
1-3-5-8-10=5+4+6+4+5=24
1-3-6-9-10=5+4+4+7+5=25
1-4-7-9-10=5+2+2+7+5=21

So the critical path = 1-3-6-9-10 as it has longest duration

Answer b=2

Number of weeks = length of critical path =25

Answer c= The options are

Cut activity 3 by 3 weeks.

and then reduce activity 6 or 9 by one week.

Reason- When activity 3 is reduced by 3 weeks, no change will occur to the critical path

In the case of reducing activity 6 or 9 by one week, it will change the critical path to  1-3-5-8-10